The small island nation of Kaboom is a simple economy with no​ government, no​ taxes, and no imports or exports. Kaboomers​ (citizens of​ Kaboom) are creatures of habit. They have a rule that everyone saves exactly 20 percent of income. Assume that planned investment is fixed and remains at 225 nmillion Kaboomian dollars. Further assume that autonomous consumption​ (independent of​ Y) is​ zero, so consumption​ (C) is MPC times Upper YMPC×Y.

The following data are estimates for the island of​ Kaboom:

Real GNP​ (Y): 844 million Kaboomian dollars  

Planned investment spending (I): 225 million Kaboomian dollars

Kaboom's economy will reach equilibrium when its real GNP​ = ______ million Kaboomian dollars.​ (Enter your response as an integer.​)
Answer #1

MPC = 1 - 0.2 = 0.8

I = 225 mn Kaboomian dollars

C = 0.8 x Y

Y = C + I (At equilibrium)

Y = 0.8Y + 225

Y = 1125 mn Kaboomian dollars